OpenStudy (anonymous):

4(3 - 2x) = 15 solve for x and leave in an improper fraction.

OpenStudy (saifoo.khan):

\[x = -\frac38\]

OpenStudy (lilg132):

12 - 8x = 15 8x = 15 - 12 8x = 3 x = 3/8

OpenStudy (lilg132):

how do you write fractions i cant find it saifoo khan

OpenStudy (anonymous):

THank you and go to equations

OpenStudy (lilg132):

where in equations

OpenStudy (lilg132):

sorry answer is x = -3/8

OpenStudy (anonymous):

idk... i tried and yeah i know.. :D

OpenStudy (lilg132):

dont forget to click good answer :p

OpenStudy (anonymous):

hey why does it end up a negative?